OpenStudy (firejay5):

0.25y = 10 in standard form

OpenStudy (anonymous):

standard form is bring evyerthing to one side. For example: y= x In standard form: y-x = 0

OpenStudy (firejay5):

stardard form is: Ax + By = 0

OpenStudy (anonymous):

y = 1y

OpenStudy (anonymous):

It's the same thing.

OpenStudy (firejay5):

yes so

OpenStudy (firejay5):

How should it be worked

OpenStudy (anonymous):

Substract 10 from both sides.

OpenStudy (anonymous):

0.25y-10 = 0 That's standard form and if you want also, you can make it a y by multiplying it by 4. y-40 = 0

OpenStudy (firejay5):

you're good at this what's your facebook??

OpenStudy (anonymous):

Just add me as a fan, you'll see when I'm online and you can ask me questions.

OpenStudy (firejay5):

aww poopy I wanted to stay in contact

OpenStudy (anonymous):

Don't worry, I'm here everyday :)

OpenStudy (firejay5):

promises

OpenStudy (firejay5):

you're so cool